### v1

# Fast and accurate emulation of complex dynamic simulators Junoh Heo

Abstract
While dynamic simulators, which are computational models that evolve over time and are governed by differential equations, are essential in scientific and engineering applications, their emulation remains challenging due to the unpredictable behavior of complex systems. To address this challenge, this paper introduces a fast and accurate Gaussian Process (GP)-based emulation method for complex dynamic simulators. By integrating linked GPs into the one-step-ahead emulation framework, the proposed algorithm enables exact analytical computations of the posterior mean and variance, eliminating the need for computationally expensive Monte Carlo approximations. This significantly reduces computation time while maintaining or improving predictive accuracy. Furthermore, the method extends naturally to systems with forcing inputs by incorporating them as additional variables within the GP framework. Numerical experiments on the Lotka-Volterra model and the Lorenz system demonstrate the efficiency and computational advantages of the proposed approach. An R package, dynemu, implementing the one-step-ahead emulation approach, is available on CRAN.
Fast and accurate emulation of complex dynamic simulators
Junoh Heo
Michigan State University
Keywords: Surrogate model, Linked Gaussian processes, Complex system, Uncertainty propagation
1 Introduction
Computer models, often referred to as computer simulators, are indispensable tools for simulating complex systems and processes to gain insights into physical or scientific phenomena. Many of these models evolve over time and are governed by intricate systems of differential equations, collectively known as dynamic simulators (Scheinerman,, 2012). Dynamic simulators are used to describe a wide range of phenomena across diverse fields, including meteorology (Millán et al.,, 2009), hydrology (Li et al.,, 2016), physics (Bahamonde et al.,, 2018), engineering (Bongard and Lipson,, 2007), and biology (Hwang et al.,, 2025), among others. These models are particularly crucial for studying problems where analytical solutions to the governing equations are either impractical or infeasible.
Dynamic systems often have forcing inputs, which are external or time-varying variables that influence the system’s evolution. Examples include external forces in mechanical systems, atmospheric conditions in climate models, or control inputs in engineering applications. These inputs are typically included as additional input variables in the emulator and are usually assumed to be known. Their presence modifies the system dynamics, necessitating an adaptation of the emulation approach to capture their effects properly.
Despite their utility, dynamic simulators often come with significant computational demands, which can limit their practicality in certain applications. These models typically require solving systems of ordinary differential equations (ODEs), which can be computationally intensive, especially when simulations must be repeated for tasks such as optimization, sensitivity analysis, or uncertainty quantification. To overcome these limitations, emulation offers an efficient alternative by constructing surrogate models that approximate the behavior of the simulator. Surrogate models emulate the simulator’s outputs at a fraction of the computational cost, enabling researchers to conduct extensive analyses without the prohibitive expense of repeated direct simulations. This capability enhances the utility of dynamic simulators, facilitating faster and more scalable solutions to complex problems across scientific and engineering domains.
GPs (Rasmussen and Williams,, 2006) are a widely used choice for constructing emulators due to their flexibility, accuracy, and strong theoretical foundation. They are particularly well-suited for approximating the complex, non-linear relationships commonly encountered in real-world problems. A key advantage of GPs lies in their ability not only to provide mean predictions but also to quantify uncertainty through analytical expressions of posteriors. Moreover, GPs naturally interpolate training data, ensuring exact predictions at observed points. This interpolation property is particularly important for dynamic simulators, which are inherently deterministic, as it preserves the consistency of the emulator with the original simulator outputs.
Several GP emulators for dynamic simulators have been developed, ranging from classic methods to recent advancements (Girard et al.,, 2002; Bhattacharya,, 2007; Conti et al.,, 2009; Mohammadi et al.,, 2024), achieving significant performance improvements. For a detailed overview of dynamic simulator emulation, we refer readers to Mohammadi et al., (2024). Notably, Bhattacharya, (2007) proposed a one-step-ahead emulation approach, which assumes that the model output at time $t$ depends only on the output at the previous time point $t-1$, a property commonly referred to as the Markov property. Building on this, Mohammadi et al., (2019) enhanced the approach by incorporating input uncertainty and emulating the numerical flow map. Despite these advancements, many emulators still require intensive computations, such as Monte Carlo (MC) approximations (Conti et al.,, 2009; Mohammadi et al.,, 2019) or grid-based methods (Bhattacharya,, 2007).
To the best of our knowledge, the posterior distribution of GPs with uncertain inputs is known to be intractable, necessitating the use of MC methods (Girard et al.,, 2002; Mohammadi et al.,, 2019) in the context of dynamic simulators. To address this limitation, Mohammadi et al., (2024) recently employed random Fourier features to approximate the kernel. However, maintaining the quality of the approximated kernel requires drawing hundreds of random features. Furthermore, generating predictions involves drawing hundreds of additional sample paths from the emulated flow map. These computational demands may limit the practical advantages of emulators, highlighting the need for more efficient approaches.
In this article, we propose a fast and more accurate emulation method for dynamic simulators by treating the emulator as self-coupled, adopting the one-step-ahead approach while overcoming the aforementioned limitations of MC approximations. The proposed method builds on the framework of Mohammadi et al., (2019) but replaces the approximated distribution with the exact mean and variance. These analytically tractable predictive posterior mean and variance, enabled by Linked GPs (Kyzyurova et al.,, 2018; Ming and Guillas,, 2021), facilitate the computation of exact posteriors, thus eliminating the need for time-consuming MC approximations. Linked GPs, originally developed for coupled computer models, allow for the analytical derivation of the mean and variance of GPs at random inputs. These closed-form expressions are achievable with commonly used kernels, such as the squared exponential kernel or the Matérn kernel (Stein,, 1999), with smoothness parameters $\nu=1.5$ and $\nu=2.5$. By leveraging this framework, the proposed model achieves computational efficiency without any loss of information, as the computational burdens of most existing methods primarily arise from the approximation step.
The remainder of this article is organized as follows. Section 2 provides a review of GP emulators, linked GPs, and the one-step-ahead approach. The proposed fast and accurate emulation method is introduced in Section 3. Section 4, presents several numerical studies to demonstrate the competitiveness of the proposed method. Finally, Section 5 concludes the paper.
2 Review
2.1 Gaussian processes
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ represent the computer code output as a deterministic black-box function of the input $\mathbf{x}$, where $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{d}$, and the corresponding output is $y=f(\mathbf{x})$. Typically, a GP prior assumes that the simulation outputs $\mathbf{y}$ follow a multivariate normal distribution:
$\mathbf{y}\sim\mathcal{N}\left(\alpha(\mathbf{x}),\tau^{2}K(\mathbf{x},\mathbf{x}^{\prime})\right),$
where $\alpha(\mathbf{x})$ is the mean function, $\tau^{2}$ is a scale parameter, and $K(\mathbf{x},\mathbf{x}^{\prime})$ is a positive-definite kernel matrix. Popular choices for the mean function $\alpha(\mathbf{x})$ include linear $(1,\mathbf{x}^{T})\beta$, constant $\alpha$, or zero, where $\beta$ is a vector of $d+1$ regression coefficients. For the kernel $K$, the squared exponential kernel or Matérn kernel is commonly adopted. In this paper, we focus on the anisotropic squared exponential kernel $K(\mathbf{x},\mathbf{x}^{\prime})=\exp{\left(-\sum^{d}_{k=1}\frac{\left(x_{k}-x_{k}^{\prime}\right)^{2}}{\theta_{k}}\right)}$, where $(\theta_{1},\ldots,\theta_{d})$ are the lengthscale hyperparameters that represent the rate of decay of correlation in each input dimension.
The posterior prediction of GPs at a new input point $\mathbf{x}$, given the input design $X=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})\subset\mathbb{R}^{n\times d}$ and corresponding outputs $\mathbf{y}$, is given as:
$\mu(\mathbf{x})$ $=\alpha(\mathbf{x})+\mathbf{k}^{T}(\mathbf{x})\mathbf{K}^{-1}\left(\mathbf{y}-\alpha(X)\right),\quad\text{and}$
$\sigma^{2}(\mathbf{x})$ $=\tau^{2}\left(1-\mathbf{k}^{T}(\mathbf{x})\mathbf{K}^{-1}\mathbf{k}(\mathbf{x})\right),$
where $\mathbf{k}(\mathbf{x})_{i}=K(X,\mathbf{x})_{i}$ and $\mathbf{K}_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})$. For more details about GPs in the context of computer experiments, we refer to Santner et al., (2018) and Gramacy, (2020).
The parameters $\left\{\beta,\tau^{2},\theta_{1},\ldots,\theta_{d}\right\}$ in the GP model can be estimated by maximizing the log-likelihood:
$-\frac{1}{2}\left(n\log 2\pi+n\log\tau^{2}+\log|\mathbf{K}|+\frac{1}{\tau^{2}}\left(\mathbf{y}-\alpha(X)\right)^{T}\mathbf{K}^{-1}\left(\mathbf{y}-\alpha(X)\right)\right).$
The profile log-likelihood is obtained by substituting the analytical expressions of $\hat{\beta}$ and $\hat{\tau}^{2}$:
$\hat{\beta}=\left(X_{n}^{T}\mathbf{K}^{-1}X_{n}\right)^{-1}X_{n}^{T}\mathbf{K}^{-1}\mathbf{y}\quad\text{and}\quad\hat{\tau}^{2}=\frac{\left(\mathbf{y}-(1,X_{n}^{T})\hat{\beta}\right)^{T}\mathbf{K}^{-1}\left(\mathbf{y}-(1,X_{n}^{T})\hat{\beta}\right)}{n}.$
In practice, optimization algorithms such as L-BFGS-B (Byrd et al.,, 1995; Zhu et al.,, 1997) are commonly employed to maximize the log-likelihood.
2.2 Linked GPs
Linked GP emulators (Kyzyurova et al.,, 2018; Ming and Guillas,, 2021) are designed to emulate two or more coupled systems of computer models. Consider $n_{1}$ computer models at the first layer, $f_{l}:\mathbb{R}^{d_{1}}\rightarrow\mathbb{R}$ for $l=1,\ldots,n_{1}$, which are connected to a second-layer model $g:\mathbb{R}^{d_{2}}\rightarrow\mathbb{R}$. In this framework, the outputs of $f_{1},\ldots,f_{n_{1}}$ serve as part of the input to $g$. Assuming that $f_{1},\ldots,f_{n_{1}}$ are independent and share the same set of inputs, the posterior predictive distribution at a new input $\mathbf{x}$, given the input-output pairs $\left(X_{n},f_{l}(X_{n})\right)_{l=1,\ldots,n_{1}}$ and $\left(Z_{n},g(Z_{n})\right)$ is expressed as:
$p$ $(g\circ(f_{1},\ldots,f_{n_{1}})(\mathbf{x})|\left\{f_{l}(X_{n})\right\}_{l=1,\ldots,n_{1}},g(Z_{n}),\mathbf{x})$
$=\int p(g\left(\mathbf{x}_{g},f_{1}(\mathbf{x}_{f}),\ldots,f_{n_{1}}(\mathbf{x}_{f})\right)|\left\{f_{l}(X_{n})\right\}_{l=1,\ldots,n_{1}},g(Z_{n}),\mathbf{x})$
×$\times\prod^{n_{1}}_{l=1}p(f_{l}(\mathbf{x}_{f})|f_{l}(X_{l}))df_{1}(\mathbf{x}_{f})\ldots df_{n_{1}}(\mathbf{x}_{f})$ (1)
where $\mathbf{x}_{g}$ denotes the unlinked inputs of $g$, and $\mathbf{x}_{f}$ represents the inputs for $f_{1},\ldots,f_{n_{1}}$, which are a subset of $\mathbf{x}$. This posterior is known to be analytically intractable (Girard et al.,, 2002). Numerical techniques, such as Markov Chain Monte Carlo, can approximate the posterior but are computationally demanding. To address this limitation, subsequent works replaced $f_{1},\ldots,f_{n_{1}}$ and $g$ with their respective GP emulators $\hat{f}_{1},\ldots,\hat{f}_{n_{1}}$ and $\hat{g}$ and applied Gaussian approximations. Using this approach, Kyzyurova et al., (2018) derived the closed-form expressions under the squared exponential kernel, and Ming and Guillas, (2021) extended the posterior to the Matérn kernels.
Linked GP emulators minimize the Kullback-Leibler divergence between the emulator and a Gaussian density (Ming and Guillas,, 2021). Moreover, they demonstrate superior performance compared to composite emulator, which disregard the coupled relationships and consider only the inputs of the first layer and the outputs of the final layer. This concept of linked GPs can be naturally extended to systems with multiple layers involving iterative or parallel structures (Ming and Guillas,, 2021; Heo and Sung,, 2023), high-dimensional output (Dolski et al.,, 2024), as well as deep GPs (Ming et al.,, 2023).
2.3 One-step-ahead emulations
Consider a dynamic computer simulator without forcing input $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times T}$ that produces $T$ outputs evolving over time. The simulator $f$ is governed by $d$ ODEs, where the state variables at time $t_{s}$ are expressed as $\mathbf{x}(t_{s})=(x_{1}(t_{s}),\ldots,x_{d}(t_{s}))^{T}$, for $s=0,\ldots,T$. The one-step-ahead approach emulates the system’s flow map over short time intervals, assuming the Markov property, which states that the state $\mathbf{x}(t_{s})$ depends only on the previous state $\mathbf{x}(t_{s-1})$. Once an emulator $\hat{f}$ is constructed based on the input-output pairs at the initial condition $\mathbf{x}(t_{0})$ and the corresponding output $\mathbf{x}(t_{1})$, predictions for future states $\hat{\mathbf{x}}(t_{s})$ for $s=1,\ldots,T$ can be obtained iteratively by leveraging the Markov assumption (Mohammadi et al.,, 2019, 2024):
$\hat{x}_{m}(t_{s+1})=\hat{f}_{m}(\mathbf{x}(t_{s})),$ (2)
where $m=1,\ldots,d$. This approach is particularly effective for periodic systems, as it can accurately predict the next state based on the current state within the input range, making it ideal for capturing the system’s repetitive dynamics. Furthermore, the one-step-ahead approach provides an efficient, reliable, and computationally inexpensive alternative to other emulation methods for dynamic simulators (Conti et al.,, 2009; Stolfi and Castiglione,, 2021).
Mohammadi et al., (2019) incorporate input uncertainty at each time step into the emulators by adopting MC methods and leveraging the laws of total expectation and total variance to approximate the posterior in (2.2):
$\mathbb{E}\left[\hat{x}_{m}(t_{s})\right]$ $=\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))\right]=\mathbb{E}\left[\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))|\mathbf{x}(t_{s-1})\right]\right],$ (3)
$\mathbb{V}\left[\hat{x}_{m}(t_{s})\right]$ $=\mathbb{V}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))\right]$
$=\mathbb{E}\left[\mathbb{V}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))|\mathbf{x}(t_{s-1})\right]\right]+\mathbb{V}\left[\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))|\mathbf{x}(t_{s-1})\right]\right].$ (4)
Modeling GPs for $f_{l}$ This framework enables the propagation of uncertainties inherent in the inputs through the emulator. Furthermore, the approach has been extended to account for correlated emulators, enabling a comprehensive treatment of dependencies across time steps.
3 Fast and exact emulation
In contrast to previously established findings, the prediction of GP with uncertain inputs in the one-step-ahead approach can be made computationally tractable by integrating the concept of Linked GP emulators. To be specific, the proposed approach predicts each state by linking GP outputs recursively, preserving accuracy without the need for sampling. Interestingly, the model structure depicted in Figure 1 can be interpreted as a special case within the emerging field of Deep GP (DGP) models (Damianou and Lawrence,, 2013), which have gained significant attention in the computer experiments literature for their hierarchical representation and enhanced flexibility (Sauer et al.,, 2023; Ming et al.,, 2023). Unlike canonical DGPs that rely on latent variable hierarchies, the proposed framework employs GP emulators fitted to each state variable, structured in a feed-forward manner reminiscent of deep learning models. This formulation facilitates the propagation of uncertainty across layers while maintaining analytical tractability, thereby enhancing predictive performance in dynamic system emulation. Building on this integration, we introduce a fast and accurate prediction algorithm for the one-step-ahead approach.
[…]
Figure 1: An illustration of the self-linked GP emulator for a dynamic system governed by $d$ ODEs. Starting with the initial state $\mathbf{x}(t_{0})$, each emulator $\hat{f}_{m}$, $m=1,\ldots,d$ predicts the $m$-th component of the state $\hat{x}_{m}(t_{s})$ for $s=1,\ldots,T$. The predicted state $\hat{\mathbf{x}}(t_{s})$ is then fed forward as the input for the next time step $s+1$.
3.1 Algorithm without forcing inputs
For a nonlinear dynamic simulator $f$ governed by $d$ ODEs, consider GP emulators $\hat{f}_{m}$, $m=1,\ldots,d$ constructed by using an initial design $\mathbf{X}=(\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{n}(t_{0}))$, and corresponding outputs $\mathbf{y}_{m}=(x_{m}^{1}(t_{1}),\ldots,x_{m}^{n}(t_{1}))^{T}$ following the one-step-ahead approach. Instead of relying on MC approximations to compute (3) and (2.3), we derive analytical expressions for $\mathbb{E}\left[\hat{x}_{m}(t_{s})\right]$ and $\mathbb{V}\left[\hat{x}_{m}(t_{s})\right]$ under the squared exponential kernel given $\hat{\mathbf{x}}(t_{s-1})$ are as follows:
$\mu_{m}(\hat{\mathbf{x}}(t_{s}))$ $=\mathbb{E}\left[\hat{x}_{m}(t_{s})\right]=\mathbb{E}\left[\hat{f}_{m}(\hat{\mathbf{x}}(t_{s-1}))\right]$
$=\begin{pmatrix}1&\hat{\mathbf{x}}(t_{s-1})\end{pmatrix}\beta_{m}+\sum^{n}_{i=1}r_{i}\prod_{l=1}^{d}\sqrt{\frac{\theta_{ml}}{\theta_{ml}+2\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}}\exp{\left(-\frac{\left(x^{i}_{l}(t_{0})-\mu_{l}(\hat{\mathbf{x}}(t_{s-1}))\right)^{2}}{\theta_{ml}+2\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}\right)},$ (5)
and
$\sigma^{2}_{m}(\hat{\mathbf{x}}(t_{s}))$ $=\mathbb{V}\left[\hat{x}_{m}(t_{s})\right]=\mathbb{V}\left[\hat{f}_{m}(\hat{\mathbf{x}}(t_{s-1}))\right]$
$=\tau^{2}_{m}-\left(\mu_{m}(\hat{\mathbf{x}}(t_{s}))-\begin{pmatrix}1&\hat{\mathbf{x}}(t_{s-1})\end{pmatrix}\beta_{m}\right)^{2}+\left(\sum_{i,k=1}^{n}\left(r_{i}r_{k}-\tau^{2}_{m}(\mathbf{K}^{-1}_{m})_{ik}\right)\prod_{l=1}^{d}\zeta_{lik}\right),$ (6)
where $r_{i}=(\mathbf{K}^{-1}_{m}(\mathbf{y}_{m}-\mathbf{X}\beta_{m}))_{i}$, and
$\zeta_{lik}=\sqrt{\frac{\theta_{ml}}{\theta_{ml}+4\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}}\exp{\left(-\frac{\left(\frac{x^{i}_{l}(t_{0})+x^{k}_{l}(t_{0})}{2}-\mu_{l}(\hat{\mathbf{x}}(t_{s-1}))\right)^{2}}{\frac{\theta_{ml}}{2}+2\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}-\frac{(x^{i}_{l}(t_{0})-x^{k}_{l}(t_{0}))^{2}}{2\theta_{ml}}\right)}.$
The closed-form expressions for the posterior predictive mean (3) and variance (2.3) under a Matérn kernel with smoothness parameters $\nu=1.5$ and $\nu=2.5$ can be straightforwardly derived following the proof in Ming and Guillas, (2021). Consistent with prior studies (Girard et al.,, 2002; Kyzyurova et al.,, 2018; Mohammadi et al.,, 2019; Ming and Guillas,, 2021), we adopt the moment matching method to approximate the posterior distribution as Gaussian, leveraging the mean and variance from (5) and (6). The moment matching approximation suits the structure of self-linked GP emulator, where outputs recursively serve as inputs, since the method aligns with the inherent Gaussian nature of GPs and ensures analytical tractability and recursive consistency.
Algorithm 1 Exact emulation of dynamic simulators without forcing inputs
1:for $m=1$ to $d$ do
2:     Fit $\hat{f}_{m}$ using $\mathbf{X}=(\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{n}(t_{0}))$, and $\mathbf{y}_{m}=(x^{1}_{m}(t_{1}),\ldots,x^{n}_{m}(t_{1}))^{T}$
3:end for
4:Set $\hat{\mathbf{x}}(t_{0})\leftarrow\mathbf{x}(t_{0})$
5:for $s=0$ to $T-1$ do
6:     for $m=1$ to $d$ do
7:         Compute $\mu_{m}(\hat{\mathbf{x}}(t_{s+1}))$ using (5)
8:         Compute $\sigma^{2}_{m}(\hat{\mathbf{x}}(t_{s+1}))$ using (6)
9:     end for
10:     Update $s\leftarrow s+1$
11:end for
12:Return: $\hat{\mathbf{x}}(t_{s})$ for $s=1,\ldots,T$
By leveraging these closed-form expressions, computationally expensive MC approximations can be avoided, allowing for exact computations. The elimination of MC sampling significantly reduces computational costs, making the proposed approach particularly suitable for dynamical systems with many state variables that require extensive simulations. Moreover, the proposed framework provides exact predictive moments, ensuring both efficiency and accuracy. This is particularly advantageous for long-term iterative predictions, where even small numerical errors in early steps can accumulate over time, leading to significant deviations in later steps—a phenomenon often referred to as the butterfly effect. By directly computing the mean and variance, our approach mitigates this issue, enhancing robustness in sequential simulations. Furthermore, multi-step-ahead predictions can be efficiently obtained by recursively applying (5) and (6), reducing the need for costly recomputation at each iteration.
3.2 Algorithm with forcing input
Assume we have a known, $d_{w}$-dimensional forcing input $\mathbf{w}(t_{s})$ at time $t_{s}$ for $s=1,\ldots,T$. The dynamic simulator $f$ can be expressed as:
$x_{m}(t_{s+1})=f_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s})),\quad\text{for}\quad m=1,\ldots,d.$ (7)
Accordingly, the Markov property must be adjusted to account for forcing inputs. Instead of (2), the emulator follows:
$\hat{x}_{m}(t_{s+1})=\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s})),$ (8)
where $m=1,\ldots,d$. The algorithm proposed in Section 3.1 can be naturally extended to incorporate forcing inputs by treating them as additional variables in the GP framework. This extension preserves the computational efficiency of the method while enabling accurate predictions for systems influenced by external inputs.
To account for forcing inputs, we modify the analytical expressions for $\mathbb{E}\left[\hat{x}_{m}(t_{s})\right]$ and $\mathbb{V}\left[\hat{x}_{m}(t_{s})\right]$ under the squared exponential kernel given $\hat{\mathbf{x}}(t_{s-1})$ are as follows:
$\mu_{m}(\hat{\mathbf{x}}(t_{s}))=\begin{pmatrix}1&\hat{\mathbf{x}}(t_{s-1})\end{pmatrix}\beta_{m}$
$+\sum^{n}_{i=1}r_{i}\prod_{j=1}^{d_{w}}\exp\left(-\frac{\left(w^{i}_{j}(t_{0})-w_{j}(t_{s})\right)^{2}}{\theta_{wj}}\right)\prod_{l=1}^{d}\frac{1}{\sqrt{1+2\frac{\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}{\theta_{ml}}}}\exp{\left(-\frac{\left(x^{i}_{l}(t_{0})-\mu_{l}(\hat{\mathbf{x}}(t_{s-1}))\right)^{2}}{\theta_{ml}+2\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}\right)},$ (9)
and
$\sigma^{2}_{m}(\hat{\mathbf{x}}(t_{s}))=\tau^{2}_{m}-\left(\mu_{m}(\hat{\mathbf{x}}(t_{s}))-\begin{pmatrix}1&\hat{\mathbf{x}}(t_{s-1})\end{pmatrix}\beta_{m}\right)^{2}$
$+\left(\sum_{i,k=1}^{n}\left(r_{i}r_{k}-\tau^{2}_{m}(\mathbf{K}^{-1}_{m})_{ik}\right)\prod_{j=1}^{d_{w}}\exp{\left(-\frac{\left(w^{i}_{j}(t_{0})-(w_{j}(t_{s})\right)^{2}+\left(w^{k}_{j}(t_{0})-w_{j}(t_{s})\right)^{2}}{\theta_{wj}}\right)}\prod_{l=1}^{d}\zeta_{lik}\right),$ (10)
where $\theta_{wj}$ for $j=1,\ldots,d_{w}$ are the lengthscale hyperparameters associated with the forcing inputs.
Algorithm 2 Exact emulation of dynamic simulators with forcing inputs
1:for $m=1$ to $d$ do
2:     Fit $\hat{f}_{m}$ using $\mathbf{X}=(\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{n}(t_{0}))$, $\mathbf{W}=(\mathbf{w}^{1}(t_{0}),\ldots,\mathbf{w}^{n}(t_{0}))$, and $\mathbf{y}_{m}=(x^{1}_{m}(t_{1}),\ldots,x^{n}_{m}(t_{1}))^{T}$
3:end for
4:Set $\hat{\mathbf{x}}(t_{0})\leftarrow\mathbf{x}(t_{0})$
5:for $s=0$ to $T-1$ do
6:     for $m=1$ to $d$ do
7:         Compute $\mu_{m}(\hat{\mathbf{x}}(t_{s+1}))$ using (9) with $\mathbf{w}(t_{s})$
8:         Compute $\sigma^{2}_{m}(\hat{\mathbf{x}}(t_{s+1}))$ using (10) with $\mathbf{w}(t_{s})$
9:     end for
10:     Update $s\leftarrow s+1$
11:end for
12:Return: $\hat{\mathbf{x}}(t_{s})$ for $s=1,\ldots,T$
The additional term in these expressions corresponds to the kernel function $K(\mathbf{w}(t_{0}),\mathbf{w}(t_{s}))$, which arises because the forcing inputs $\mathbf{w}(t_{s})$ are treated as known rather than random. This assumption simplifies the model structure, as the forcing inputs remain unlinked variables in the emulator.
4 Numerical Studies
In this section, we conduct a series of numerical experiments to evaluate the performance of the proposed approach. Specifically, we compare the predictive performance of the proposed exact algorithm with an emulation method based on MC approximation, using $n_{MC}=1000$ samples. Both approaches employ the same GP emulator, using an anisotropic squared exponential kernel and a linear mean function $\alpha(\mathbf{x})=(1,\mathbf{x}^{T})\beta$.
All numerical studies are conducted using a maximin Latin hypercube design (Johnson et al.,, 1990; Morris and Mitchell,, 1995), with the number of initial samples set to $n=12d$, following the guidelines of Mohammadi et al., (2019), which builds on the design principles established in Jones et al., (1998) and Loeppky et al., (2009). The time step difference is set to $\Delta t=t_{s}-t_{s-1}=0.01$.
The governing ODEs of the dynamic simulators are solved using the lsoda method from the R package deSolve (Soetaert et al.,, 2010). This solver, known as the Livermore solver for stiff and non-stiff ODE systems, automatically switches between stiff and non-stiff methods for efficiency (Petzold,, 1983).
To evaluate predictive performance, we compute predictions at all time points $t_{s}$, $s=1,\ldots,T$, and assess model accuracy using the root-mean-square error (RMSE) and continuous rank probability score (CRPS) (Gneiting and Raftery,, 2007) over 100 repetitions. The evaluation metrics are defined as follows:
RMSE $=\sqrt{\sum_{i=1}^{n}\frac{\left(y_{i}-\hat{y}_{i}\right)^{2}}{n}},$
CRPS $=\sigma\left[\frac{1}{\sqrt{\pi}}-2\phi\left(\frac{y-\mu}{\sigma}\right)-\frac{y-\mu}{\sigma}\left(2\Phi\left(\frac{y-\mu}{\sigma}\right)-1\right)\right],$
where $y$ represents the true function values, $\mu$ and $\sigma^{2}$ denote the posterior predictive mean and variance, respectively. $\phi$ and $\Phi$ are the pdf and the cdf of a standard Gaussian distribution. These metrics quantify the degree of discrepancy between the true function and the predictive model, with lower RMSE and CRPS values indicating better model accuracy. Furthermore, we evaluate computational efficiency by comparing computation times across different emulation approaches.
4.1 Lotka-Volterra model
The Lotka-Volterra model (Lotka,, 1925; Volterra,, 1927; Goel et al.,, 1971), also known as the consumer-prey model, is a foundational framework in mathematical ecology for describing predator-prey dynamics. This model characterizes the temporal evolution of two interacting populations: a prey species with population density $P(t)$ and a predator species with population density $C(t)$. The dynamics are governed by the following system of nonlinear differential equations:
$\begin{cases}\frac{dP}{dt}&=r_{G}\cdot P\cdot\left(1-\frac{P}{K}\right)-r_{I}\cdot P\cdot C,\\ \frac{dC}{dt}&=k_{AE}\cdot r_{I}\cdot P\cdot C-r_{M}\cdot C,\end{cases}$
where $r_{G}>0$ represents the intrinsic growth rate of the prey in the absence of predators, $K$ is the prey carrying capacity, $r_{I}>0$ is the predation rate coefficient, $k_{AE}>0$ is the assimilation efficiency of the predator, and $r_{M}>0$ denotes the natural mortality rate of the predator in the absence of prey. For this study, we consider a simulation over $t\in[0,30]$, with prespecified parameter values $r_{G}=1.5$, $K=10$, $r_{I}=1.5$, $k_{AE}=1$, and $r_{M}=2$. Initial samples are drawn from $[0,5]^{2}$. The objective is to predict the system’s dynamics given the initial state $\mathbf{x}(t_{0})=(P(t_{0})=1,C(t_{0})=2)$.
Figure 2: RMSE (first and second plots), CRPS (third and fourth plots) of emulation algorithms to $P$ and $C$ and computation time (fifth plots) across 100 repetitions for Lotka-Volterra model.
Figure 3: Lotka-Volterra model (black solid line) and emulation results of prediction (green dashed line) via each algorithm.
Figure 2 presents the emulation results over 100 repetitions for RMSE and CRPS, while Figure 3 illustrates the predicted trajectories for state variables, $P$ and $C$, respectively. As expected, both algorithms produce nearly identical predictions, with a maximum difference of only 0.000024. While both methods achieve high predictive accuracy, their computational costs differ considerably. On average, the exact algorithm requires only 0.40603 seconds per repetition, whereas the MC approach takes 90.13912 seconds. This comparison highlights the advantage of the proposed framework in significantly reducing computational cost while maintaining accuracy in dynamic system emulation. Moreover, the computational gap would increase further if the MC approach employed a larger number of samples, or if the simulation involved many time steps or a longer time horizon.
Figure 4: Boxplots of computation time for the Exact and Monte Carlo methods as the number of design points varies from 5 to 50 in increments of 5 across 100 repetitions for Lotka-Volterra model.
Figure 4 illustrates the computation time for emulation as the number of samples varies from 5 to 50 per repetition. With a precomputed $\mathbf{K}^{-1}$, the computational complexity of computing the predictive mean and variance at a single input location using MC-based emulation is $\mathcal{O}(n^{2}n_{MC})$. On the other hand, under the same conditions, the exact algorithm requires only $\mathcal{O}(n^{2}d)$ complexity. Since $n_{MC}$ is typically in the range of several thousand or more, the proposed algorithm offers substantial computational savings.
Figure 5: RMSE (first plot), CRPS (second plot) of emulation algorithms to $P$ with forcing input $C$ and computation time (third plot) across 100 repetitions for Lotka-Volterra model.
Figure 5 demonstrates the results for the case in which $C(t)$ is treated as a forcing input. As anticipated, the emulation performance improves due to the availability of a known input, and the computation time is halved with 0.20403 seconds per repetition for the exact algorithm and 48.03827 seconds for the MC-based algorithm. These results confirm that the exact method remains computationally efficient in the presence of a forcing input.
4.2 Lorenz system
The Lorenz system, introduced by Lorenz, (1963), models the behavior of atmospheric convection through a set of three differential equations:
$\begin{cases}\frac{dX}{dt}&=aX+YZ,\\ \frac{dY}{dt}&=b(Y-Z),\\ \frac{dZ}{dt}&=-XY+cY-Z,\end{cases}$
where $X$, $Y$, and $Z$ represent the system’s state variables, and $a$, $b$, and $c$ are system parameters that govern the system’s dynamics. Using the classic values $a=-8/3$, $b=-10$, and $c=28$, originally employed by Lorenz to study convection rolls, the system exhibits chaotic behavior. We focus on a simulation over $t\in[0,25]$, and the initial state $\mathbf{x}(t_{0})=(X(t_{0})=-1,Y(t_{0})=-1,Z(t_{0})=-1)$. Initial designs are sampled from $[-10,10]^{3}$.
Figure 6: RMSE (first to third plots), CRPS (fourth to sixth plots) of emulation algorithms to $P$ and $C$ and computation time (seventh plots) across 100 repetitions for Lorenz system.
Figure 7: Lorenz system (black solid line) and emulation results of prediction (green dashed line) via each algorithm.
Figure 8: 3-d illustrations of Lorenz system (black solid line) and emulation results of prediction (green dashed line) via each algorithm.
From Figure 6, the emulation results show slight differences between the two algorithms. This discrepancy arises because the Lorenz system is highly sensitive to small changes, which can lead to significant differences, a phenomenon known as the “butterfly effect”. As shown, the exact algorithm can emulate the system efficiently, requiring only 0.95777 seconds per repetition, whereas the MC approach takes significantly longer at 143.22403 seconds. Figures 7 and 8 illustrate the predicted trajectories for each state variable, $X$, $Y$, and $Z$. Both algorithms provide accurate predictions at early time steps up to $t=20$, but their performance deteriorates beyond this point. For highly sensitive dynamic systems such as the Lorenz system, the exact algorithm is particularly advantageous, as MC methods introduce variations that exacerbate prediction errors over time. This underscores the importance of exact emulation for improving accuracy and computational efficiency in chaotic systems.
5 Discussion
In this paper, we propose fast and exact algorithms for emulating complex dynamic simulators. These algorithms enable efficient emulation compared to traditional MC-based approaches, while retaining their advantages. Numerical studies demonstrate that our method achieves equivalent results with significantly reduced computation time, over 100 times faster. Additionally, our approach accurately computes the predictive posterior mean and variance, preventing the accumulation of small differences from sampling errors.
From a computational perspective, the addition of jitter for numerical stability–commonly used to avoid ill-conditioned covariance matrices–may introduce minor discrepancies that could accumulate over time. In dynamic systems that are highly sensitive to small changes, such as those exhibiting the butterfly effect, these numerical artifacts can significantly impact the emulation outcomes. Furthermore, computational efficiency can be further improved by implementing low-level programming languages for operations.
In this work, although the state variables are correlated, we followed the approach of Damianou and Lawrence, (2013) and independently constructed the emulators. However, as suggested by Mohammadi et al., (2019), neglecting these correlations might lead to information loss, potentially affecting predictive accuracy. Future work could address this limitation by incorporating correlations among state variables, further enhancing emulation performance.
Another promising direction for future research lies in the emulation of non-periodic chaotic systems, such as the three-body problem (Newton,, 1687; Musielak and Quarles,, 2014). Chaotic systems are inherently unpredictable and highly sensitive to initial conditions, making them particularly challenging to emulate. To enhance model accuracy in such cases, integrating active learning, where new training points are adaptively selected to improve emulation performance could be a valuable approach for dynamic system emulation.

### v2

@@
-Fast and accurate emulation of complex dynamic simulators Junoh Heo
+Fast and accurate emulation of complex dynamic simulators
 Abstract
-While dynamic simulators, which are computational models that evolve over time and are governed by differential equations, are essential in scientific and engineering applications, their emulation remains challenging due to the unpredictable behavior of complex systems.
+Dynamic simulators are computational models governed by differential equations that evolve over time.
+They are essential for scientific and engineering applications but remain challenging to emulate because of the unpredictable behavior of complex systems.
 To address this challenge, this paper introduces a fast and accurate Gaussian Process (GP)-based emulation method for complex dynamic simulators.
-By integrating linked GPs into the one-step-ahead emulation framework, the proposed algorithm enables exact analytical computations of the posterior mean and variance, eliminating the need for computationally expensive Monte Carlo approximations.
-This significantly reduces computation time while maintaining or improving predictive accuracy.
-Furthermore, the method extends naturally to systems with forcing inputs by incorporating them as additional variables within the GP framework.
-Numerical experiments on the Lotka-Volterra model and the Lorenz system demonstrate the efficiency and computational advantages of the proposed approach.
-An R package, dynemu, implementing the one-step-ahead emulation approach, is available on CRAN.
+By integrating linked GPs into the one-step-ahead emulation framework, the proposed algorithm provides exact and tractable computation of the posterior mean and variance, solving a problem previously considered computationally intractable and eliminating the need for expensive Monte Carlo approximations.
+This approach substantially reduces computation time while maintaining or improving predictive accuracy.
+Furthermore, the method naturally extends to systems with forcing inputs by incorporating them as additional variables within the GP framework.
+Numerical experiments on multiple dynamic systems demonstrate the efficiency and computational advantages of the proposed approach.
+An R package, dynemu, which implements the one-step-ahead emulation approach, is available on CRAN.
 Fast and accurate emulation of complex dynamic simulators
-Junoh Heo
+Junoh Heo***Corresponding author. These authors gratefully acknowledge funding from NSF DMS 2338018.
 Michigan State University
@@
 These models are particularly crucial for studying problems where analytical solutions to the governing equations are either impractical or infeasible.
-Dynamic systems often have forcing inputs, which are external or time-varying variables that influence the system’s evolution.
-Examples include external forces in mechanical systems, atmospheric conditions in climate models, or control inputs in engineering applications.
-These inputs are typically included as additional input variables in the emulator and are usually assumed to be known.
-Their presence modifies the system dynamics, necessitating an adaptation of the emulation approach to capture their effects properly.
 Despite their utility, dynamic simulators often come with significant computational demands, which can limit their practicality in certain applications.
@@
 This capability enhances the utility of dynamic simulators, facilitating faster and more scalable solutions to complex problems across scientific and engineering domains.
-GPs (Rasmussen and Williams,, 2006) are a widely used choice for constructing emulators due to their flexibility, accuracy, and strong theoretical foundation.
-They are particularly well-suited for approximating the complex, non-linear relationships commonly encountered in real-world problems.
-A key advantage of GPs lies in their ability not only to provide mean predictions but also to quantify uncertainty through analytical expressions of posteriors.
-Moreover, GPs naturally interpolate training data, ensuring exact predictions at observed points.
-This interpolation property is particularly important for dynamic simulators, which are inherently deterministic, as it preserves the consistency of the emulator with the original simulator outputs.
-Several GP emulators for dynamic simulators have been developed, ranging from classic methods to recent advancements (Girard et al.,, 2002; Bhattacharya,, 2007; Conti et al.,, 2009; Mohammadi et al.,, 2024), achieving significant performance improvements.
+Gaussian processes (GPs) (Rasmussen and Williams,, 2006) are a widely used choice for constructing emulators due to their flexibility and strong theoretical foundation.
+They are well-suited for approximating complex, nonlinear relationships commonly encountered in real-world problems, although their performance can be sensitive to the choice of kernel and modeling assumptions.
+A key advantage of GPs lies in their ability not only to provide mean predictions but also to quantify uncertainty through analytical posterior expressions.
+In deterministic settings, GPs interpolate training data, a property particularly important for dynamic systems governed by deterministic equations, as it ensures consistency between the emulator and the original simulator outputs.
+Several GP emulators for dynamic simulators have been developed, ranging from classic methods to recent advancements (Girard et al.,, 2002; Bhattacharya,, 2007; Conti et al.,, 2009; Mohammadi et al.,, 2024).
 For a detailed overview of dynamic simulator emulation, we refer readers to Mohammadi et al., (2024).
-Notably, Bhattacharya, (2007) proposed a one-step-ahead emulation approach, which assumes that the model output at time $t$ depends only on the output at the previous time point $t-1$, a property commonly referred to as the Markov property.
+Notably, Bhattacharya, (2007) proposed a one-step-ahead emulation approach, which assumes that the model output at time $t$ depends only on the output at the previous time point t1t-1, a property commonly referred to as the Markov property.
 Building on this, Mohammadi et al., (2019) enhanced the approach by incorporating input uncertainty and emulating the numerical flow map.
@@
 These computational demands may limit the practical advantages of emulators, highlighting the need for more efficient approaches.
-In this article, we propose a fast and more accurate emulation method for dynamic simulators by treating the emulator as self-coupled, adopting the one-step-ahead approach while overcoming the aforementioned limitations of MC approximations.
-The proposed method builds on the framework of Mohammadi et al., (2019) but replaces the approximated distribution with the exact mean and variance.
-These analytically tractable predictive posterior mean and variance, enabled by Linked GPs (Kyzyurova et al.,, 2018; Ming and Guillas,, 2021), facilitate the computation of exact posteriors, thus eliminating the need for time-consuming MC approximations.
-Linked GPs, originally developed for coupled computer models, allow for the analytical derivation of the mean and variance of GPs at random inputs.
-These closed-form expressions are achievable with commonly used kernels, such as the squared exponential kernel or the Matérn kernel (Stein,, 1999), with smoothness parameters $\nu=1.5$ and $\nu=2.5$.
-By leveraging this framework, the proposed model achieves computational efficiency without any loss of information, as the computational burdens of most existing methods primarily arise from the approximation step.
+In this article, we propose a fast and accurate emulation method for dynamic simulators by treating the emulator as self-coupled and adopting a one-step-ahead approach, thereby addressing the limitations of MC approximations.
+The proposed method builds on the framework of Mohammadi et al., (2019) but replaces the MC-based estimation of the posterior with exact closed-form expressions for the predictive mean and variance.
+These analytically tractable expressions are enabled by linked GPs (Kyzyurova et al.,, 2018; Ming and Guillas,, 2021), originally developed for coupled computer models.
+In the context of dynamic emulation, recursive prediction introduces uncertainty at each step, making the inputs to the emulator uncertain inputs (i.e., stochastic inputs represented as random variables) rather than fixed values.
+The linked GP formulation allows this uncertainty to be propagated analytically over time, eliminating the need for repeated sampling and significantly reducing computational cost.
+By leveraging this framework, our method achieves computational efficiency without compromising accuracy, as most of the computational burdens in existing methods arise from the approximation step.
 The remainder of this article is organized as follows.
@@
 2.1 Gaussian processes
-Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ represent the computer code output as a deterministic black-box function of the input $\mathbf{x}$, where $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{d}$, and the corresponding output is $y=f(\mathbf{x})$.
+Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ denote a deterministic black-box function representing the output of a computer model.
+The input design matrix is given by $X=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})^{\top}\in\mathbb{R}^{n\times d}$, where each $\mathbf{x}_{i}\in\mathbb{R}^{d}$ is a design point.
+The corresponding outputs are collected in the vector $\mathbf{y}=(y_{1},\ldots,y_{n})^{\top}$ with each response defined by $y_{i}=f(\mathbf{x}_{i})$.
 Typically, a GP prior assumes that the simulation outputs $\mathbf{y}$ follow a multivariate normal distribution:
-$\mathbf{y}\sim\mathcal{N}\left(\alpha(\mathbf{x}),\tau^{2}K(\mathbf{x},\mathbf{x}^{\prime})\right),$
-where $\alpha(\mathbf{x})$ is the mean function, $\tau^{2}$ is a scale parameter, and $K(\mathbf{x},\mathbf{x}^{\prime})$ is a positive-definite kernel matrix.
-Popular choices for the mean function $\alpha(\mathbf{x})$ include linear $(1,\mathbf{x}^{T})\beta$, constant $\alpha$, or zero, where $\beta$ is a vector of $d+1$ regression coefficients.
-For the kernel $K$, the squared exponential kernel or Matérn kernel is commonly adopted.
-In this paper, we focus on the anisotropic squared exponential kernel $K(\mathbf{x},\mathbf{x}^{\prime})=\exp{\left(-\sum^{d}_{k=1}\frac{\left(x_{k}-x_{k}^{\prime}\right)^{2}}{\theta_{k}}\right)}$, where $(\theta_{1},\ldots,\theta_{d})$ are the lengthscale hyperparameters that represent the rate of decay of correlation in each input dimension.
-The posterior prediction of GPs at a new input point $\mathbf{x}$, given the input design $X=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})\subset\mathbb{R}^{n\times d}$ and corresponding outputs $\mathbf{y}$, is given as:
-$\mu(\mathbf{x})$ $=\alpha(\mathbf{x})+\mathbf{k}^{T}(\mathbf{x})\mathbf{K}^{-1}\left(\mathbf{y}-\alpha(X)\right),\quad\text{and}$
-$\sigma^{2}(\mathbf{x})$ $=\tau^{2}\left(1-\mathbf{k}^{T}(\mathbf{x})\mathbf{K}^{-1}\mathbf{k}(\mathbf{x})\right),$
-where $\mathbf{k}(\mathbf{x})_{i}=K(X,\mathbf{x})_{i}$ and $\mathbf{K}_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})$.
+$\mathbf{y}\sim\mathcal{N}_{n}\left(\alpha(\cdot),\tau^{2}K(\cdot,\cdot)\right),$
+where $\alpha(\cdot)$ is the mean function, $\tau^{2}$ is a scale parameter, and $K(\cdot,\cdot^{\prime})$ is a positive-definite kernel function.
+Popular choices for the mean function $\alpha(\mathbf{x})$ include linear $(1,\mathbf{x}^{\top})\beta$, constant $\alpha$, or zero, where $\beta$ is a vector of $d+1$ regression coefficients.
+In this paper, we adopt the linear mean function $\alpha(\mathbf{x})=(1,\mathbf{x}^{\top})\beta$.
+For the kernel $K$, the squared exponential kernel or Matérn kernel (Stein,, 1999) is commonly adopted.
+In this paper, we focus on the anisotropic squared exponential kernel $K(\mathbf{x},\mathbf{x}^{\prime})=\exp{\left(-\sum^{d}_{k=1}\frac{\left(x_{k}-x_{k}^{\prime}\right)^{2}}{\theta_{k}}\right)}$, where $\mathbf{x}=(x_{1},\ldots,x_{d})^{\top}$ and $\mathbf{x}^{\prime}=(x_{1}^{\prime},\ldots,x_{d}^{\prime})^{\top}$ are input vectors of size $d\times 1$, and $(\theta_{1},\ldots,\theta_{d})$ are the lengthscale hyperparameters representing the rate of correlation decay in each input dimension.
+The posterior prediction of GPs at a new input point $\mathbf{x}$, given the input design $X$ and corresponding outputs $\mathbf{y}$, is given as:
+$\mu(\mathbf{x})$ $=\alpha(\mathbf{x})+\mathbf{k}^{\top}(\mathbf{x})\mathbf{K}^{-1}\left(\mathbf{y}-\alpha(X)\right),\quad\text{and}$
+$\sigma^{2}(\mathbf{x})$ $=\tau^{2}\left(1-\mathbf{k}^{\top}(\mathbf{x})\mathbf{K}^{-1}\mathbf{k}(\mathbf{x})\right),$
+where $\mathbf{k}(\mathbf{x})_{i}=K(X_{i},\mathbf{x})$ and $\mathbf{K}_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})$.
 For more details about GPs in the context of computer experiments, we refer to Santner et al., (2018) and Gramacy, (2020).
-The parameters $\left\{\beta,\tau^{2},\theta_{1},\ldots,\theta_{d}\right\}$ in the GP model can be estimated by maximizing the log-likelihood:
-$-\frac{1}{2}\left(n\log 2\pi+n\log\tau^{2}+\log|\mathbf{K}|+\frac{1}{\tau^{2}}\left(\mathbf{y}-\alpha(X)\right)^{T}\mathbf{K}^{-1}\left(\mathbf{y}-\alpha(X)\right)\right).$
+The model parameters $\left\{\beta,\tau^{2},\theta_{1},\ldots,\theta_{d}\right\}$ are estimated by maximizing the log-likelihood:
+$-\frac{1}{2}\left(n\log 2\pi+n\log\tau^{2}+\log|\mathbf{K}|+\frac{1}{\tau^{2}}\left(\mathbf{y}-\alpha(X)\right)^{\top}\mathbf{K}^{-1}\left(\mathbf{y}-\alpha(X)\right)\right).$
 The profile log-likelihood is obtained by substituting the analytical expressions of $\hat{\beta}$ and $\hat{\tau}^{2}$:
-$\hat{\beta}=\left(X_{n}^{T}\mathbf{K}^{-1}X_{n}\right)^{-1}X_{n}^{T}\mathbf{K}^{-1}\mathbf{y}\quad\text{and}\quad\hat{\tau}^{2}=\frac{\left(\mathbf{y}-(1,X_{n}^{T})\hat{\beta}\right)^{T}\mathbf{K}^{-1}\left(\mathbf{y}-(1,X_{n}^{T})\hat{\beta}\right)}{n}.$
-In practice, optimization algorithms such as L-BFGS-B (Byrd et al.,, 1995; Zhu et al.,, 1997) are commonly employed to maximize the log-likelihood.
+$\hat{\beta}=\left(X_{n}^{\top}\mathbf{K}^{-1}X_{n}\right)^{-1}X_{n}^{\top}\mathbf{K}^{-1}\mathbf{y}\quad\text{and}\quad\hat{\tau}^{2}=\frac{\left(\mathbf{y}-(1,X_{n}^{\top})\hat{\beta}\right)^{\top}\mathbf{K}^{-1}\left(\mathbf{y}-(1,X_{n})\hat{\beta}\right)}{n}.$
+All hyperparameters in this paper are estimated by maximum likelihood using the L-BFGS-B algorithm (Byrd et al.,, 1995; Zhu et al.,, 1997).
 2.2 Linked GPs
 Linked GP emulators (Kyzyurova et al.,, 2018; Ming and Guillas,, 2021) are designed to emulate two or more coupled systems of computer models.
-Consider $n_{1}$ computer models at the first layer, $f_{l}:\mathbb{R}^{d_{1}}\rightarrow\mathbb{R}$ for $l=1,\ldots,n_{1}$, which are connected to a second-layer model $g:\mathbb{R}^{d_{2}}\rightarrow\mathbb{R}$.
-In this framework, the outputs of $f_{1},\ldots,f_{n_{1}}$ serve as part of the input to $g$.
-Assuming that $f_{1},\ldots,f_{n_{1}}$ are independent and share the same set of inputs, the posterior predictive distribution at a new input $\mathbf{x}$, given the input-output pairs $\left(X_{n},f_{l}(X_{n})\right)_{l=1,\ldots,n_{1}}$ and $\left(Z_{n},g(Z_{n})\right)$ is expressed as:
-$p$ $(g\circ(f_{1},\ldots,f_{n_{1}})(\mathbf{x})|\left\{f_{l}(X_{n})\right\}_{l=1,\ldots,n_{1}},g(Z_{n}),\mathbf{x})$
-$=\int p(g\left(\mathbf{x}_{g},f_{1}(\mathbf{x}_{f}),\ldots,f_{n_{1}}(\mathbf{x}_{f})\right)|\left\{f_{l}(X_{n})\right\}_{l=1,\ldots,n_{1}},g(Z_{n}),\mathbf{x})$
-×$\times\prod^{n_{1}}_{l=1}p(f_{l}(\mathbf{x}_{f})|f_{l}(X_{l}))df_{1}(\mathbf{x}_{f})\ldots df_{n_{1}}(\mathbf{x}_{f})$ (1)
-where $\mathbf{x}_{g}$ denotes the unlinked inputs of $g$, and $\mathbf{x}_{f}$ represents the inputs for $f_{1},\ldots,f_{n_{1}}$, which are a subset of $\mathbf{x}$.
+Consider $n$ computer models at the first layer, $f_{l}$ for $l=1,\ldots,n$, which are connected to a second-layer model $g$.
+In this framework, the outputs from $f_{1},\ldots,f_{n}$ serve as part of the input to $g$.
+Assuming $f_{1},\ldots,f_{n}$ are independent, the posterior predictive distribution at a new input $\mathbf{x}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n},\mathbf{x}_{g})$ where $\mathbf{x}_{l}$ is the input to $f_{l}$ and $\mathbf{x}_{g}$ is the unlinked input to $g$, is expressed as:
+$p$ $(g\circ(f_{1},\ldots,f_{n})(\mathbf{x}))=\int p(g\left(\mathbf{x}_{g},f_{1}(\mathbf{x}_{1}),\ldots,f_{n}(\mathbf{x}_{n})\right))\prod^{n}_{l=1}p(f_{l}(\mathbf{x}_{l}))df_{1}(\mathbf{x}_{f})\ldots df_{n}(\mathbf{x}_{f}).$ (1)
 This posterior is known to be analytically intractable (Girard et al.,, 2002).
 Numerical techniques, such as Markov Chain Monte Carlo, can approximate the posterior but are computationally demanding.
-To address this limitation, subsequent works replaced $f_{1},\ldots,f_{n_{1}}$ and $g$ with their respective GP emulators $\hat{f}_{1},\ldots,\hat{f}_{n_{1}}$ and $\hat{g}$ and applied Gaussian approximations.
-Using this approach, Kyzyurova et al., (2018) derived the closed-form expressions under the squared exponential kernel, and Ming and Guillas, (2021) extended the posterior to the Matérn kernels.
+To address this limitation, subsequent works replaced $f_{1},\ldots,f_{n}$ and $g$ with their respective GP emulators $\hat{f}_{1},\ldots,\hat{f}_{n}$ and $\hat{g}$ and applied Gaussian approximations.
+Using this approach, Kyzyurova et al., (2018) derived the closed-form expressions for the posterior mean and variance under the squared exponential kernel, while Ming and Guillas, (2021) extended these results to Matérn kernels.
 Linked GP emulators minimize the Kullback-Leibler divergence between the emulator and a Gaussian density (Ming and Guillas,, 2021).
-Moreover, they demonstrate superior performance compared to composite emulator, which disregard the coupled relationships and consider only the inputs of the first layer and the outputs of the final layer.
+Moreover, they demonstrate superior performance compared to the composite emulator, which disregards the coupled relationships and considers only the inputs of the first layer and the outputs of the final layer.
 This concept of linked GPs can be naturally extended to systems with multiple layers involving iterative or parallel structures (Ming and Guillas,, 2021; Heo and Sung,, 2023), high-dimensional output (Dolski et al.,, 2024), as well as deep GPs (Ming et al.,, 2023).
 2.3 One-step-ahead emulations
-Consider a dynamic computer simulator without forcing input $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times T}$ that produces $T$ outputs evolving over time.
-The simulator $f$ is governed by $d$ ODEs, where the state variables at time $t_{s}$ are expressed as $\mathbf{x}(t_{s})=(x_{1}(t_{s}),\ldots,x_{d}(t_{s}))^{T}$, for $s=0,\ldots,T$.
-The one-step-ahead approach emulates the system’s flow map over short time intervals, assuming the Markov property, which states that the state $\mathbf{x}(t_{s})$ depends only on the previous state $\mathbf{x}(t_{s-1})$.
-Once an emulator $\hat{f}$ is constructed based on the input-output pairs at the initial condition $\mathbf{x}(t_{0})$ and the corresponding output $\mathbf{x}(t_{1})$, predictions for future states $\hat{\mathbf{x}}(t_{s})$ for $s=1,\ldots,T$ can be obtained iteratively by leveraging the Markov assumption (Mohammadi et al.,, 2019, 2024):
-$\hat{x}_{m}(t_{s+1})=\hat{f}_{m}(\mathbf{x}(t_{s})),$ (2)
-where $m=1,\ldots,d$.
-This approach is particularly effective for periodic systems, as it can accurately predict the next state based on the current state within the input range, making it ideal for capturing the system’s repetitive dynamics.
+Consider a dynamic computer simulator $f$ governed by a set of $d$ ODEs, where the state at time $t_{s}$ is denoted by $\mathbf{x}(t_{s})=(x_{1}(t_{s}),\ldots,x_{d}(t_{s}))^{\top}\in\mathbb{R}^{d}$, for $s=0,\ldots,T$.
+The simulator may also include $p$-dimensional forcing inputs $\mathbf{w}(t_{s})=(w_{1}(t_{s}),\ldots,w_{p}(t_{s}))^{\top}$, which are typically assumed to be known.
+The simulator maps the current state and forcing inputs to the next state,
+$f(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))=\begin{bmatrix}f_{1}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))\\ \vdots\\ f_{d}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))\end{bmatrix}=\begin{bmatrix}x_{1}(t_{s+1})\\ \vdots\\ x_{d}(t_{s+1})\end{bmatrix}=\mathbf{x}(t_{s+1}),$
+thereby generating $d$ output trajectories over $T$ time steps.
+The one-step-ahead approach emulates the system’s flow map over short time intervals, assuming the Markov property: the next state $\mathbf{x}(t_{s+1})$ depends only on the current state $\mathbf{x}(t_{s})$ and the forcing input $\mathbf{w}(t_{s})$.
+Once a set of emulators $\{\hat{f}_{m}\}_{m=1}^{d}$ is constructed using training data $\{(\mathbf{x}^{i}(t_{0}),\mathbf{w}^{i}(t_{0})),\mathbf{x}^{i}(t_{1})\}_{i=1}^{n}$, future states $\mathbf{x}(t_{s+1})$ can be predicted recursively by leveraging the Markov assumption (Mohammadi et al.,, 2019, 2024).
+Importantly, this one-step-ahead approach requires only pairs of states one step apart rather than full trajectories, thereby substantially reducing the need to solve ODEs across all time steps.
+For notational simplicity, we write $\mathbf{x}(t_{s})$ for predicted states for $s\geq 1$.
+Under this convention, the recursive prediction is given by:
+$x_{m}(t_{s+1})=\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s})),$ (2)
+for $m=1,\ldots,d$, and $s=0,\ldots,T-1$.
+This approach is particularly effective for periodic systems, as it can accurately predict the next state based on the current state within the input range, making it well-suited for capturing repetitive dynamics.
 Furthermore, the one-step-ahead approach provides an efficient, reliable, and computationally inexpensive alternative to other emulation methods for dynamic simulators (Conti et al.,, 2009; Stolfi and Castiglione,, 2021).
-Mohammadi et al., (2019) incorporate input uncertainty at each time step into the emulators by adopting MC methods and leveraging the laws of total expectation and total variance to approximate the posterior in (2.2):
-$\mathbb{E}\left[\hat{x}_{m}(t_{s})\right]$ $=\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))\right]=\mathbb{E}\left[\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))|\mathbf{x}(t_{s-1})\right]\right],$ (3)
-$\mathbb{V}\left[\hat{x}_{m}(t_{s})\right]$ $=\mathbb{V}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))\right]$
-$=\mathbb{E}\left[\mathbb{V}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))|\mathbf{x}(t_{s-1})\right]\right]+\mathbb{V}\left[\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))|\mathbf{x}(t_{s-1})\right]\right].$ (4)
-Modeling GPs for $f_{l}$ This framework enables the propagation of uncertainties inherent in the inputs through the emulator.
-Furthermore, the approach has been extended to account for correlated emulators, enabling a comprehensive treatment of dependencies across time steps.
+Mohammadi et al., (2019) extends this framework by incorporating input uncertainty at each time step by adopting MC methods, and approximating the posterior in (1) via the laws of total expectation and total variance:
+$\mathbb{E}\left[x_{m}(t_{s+1})\right]$ $=\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))\right]=\mathbb{E}_{\mathbf{x}(t_{s})}\left[\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))|\mathbf{x}(t_{s}),\mathbf{w}(t_{s})\right]\right],$ (3)
+$\mathbb{V}\left[x_{m}(t_{s+1})\right]$ $=\mathbb{V}\left[\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))\right]$
+$=\mathbb{E}_{\mathbf{x}(t_{s})}\left[\mathbb{V}\left[\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))|\mathbf{x}(t_{s}),\mathbf{w}(t_{s})\right]\right]+\mathbb{V}_{\mathbf{x}(t_{s})}\left[\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))|\mathbf{x}(t_{s}),\mathbf{w}(t_{s})\right]\right].$ (4)
+In practice, equations (3) and (2.3) are approximated by drawing $n_{\mathrm{MC}}$ random samples of the uncertain inputs at every prediction step and evaluating the GP emulator at each sample.
+This procedure is repeated sequentially across all $T$ prediction steps, causing the total computational cost to increase with $n_{\mathrm{MC}}$ and $T$.
+While this framework effectively propagates input uncertainties through the emulator and has been extended to account for correlated emulators, Mohammadi et al., (2019) considers only systems without forcing inputs, limiting its flexibility for simulators with external covariates.
 3 Fast and exact emulation
-In contrast to previously established findings, the prediction of GP with uncertain inputs in the one-step-ahead approach can be made computationally tractable by integrating the concept of Linked GP emulators.
-To be specific, the proposed approach predicts each state by linking GP outputs recursively, preserving accuracy without the need for sampling.
-Interestingly, the model structure depicted in Figure 1 can be interpreted as a special case within the emerging field of Deep GP (DGP) models (Damianou and Lawrence,, 2013), which have gained significant attention in the computer experiments literature for their hierarchical representation and enhanced flexibility (Sauer et al.,, 2023; Ming et al.,, 2023).
-Unlike canonical DGPs that rely on latent variable hierarchies, the proposed framework employs GP emulators fitted to each state variable, structured in a feed-forward manner reminiscent of deep learning models.
-This formulation facilitates the propagation of uncertainty across layers while maintaining analytical tractability, thereby enhancing predictive performance in dynamic system emulation.
-Building on this integration, we introduce a fast and accurate prediction algorithm for the one-step-ahead approach.
+The one-step-ahead emulation framework can be interpreted as a probabilistic analog of the Euler method (Biswas et al.,, 2013) for solving ODEs.
+The classical Euler scheme approximates the evolution of a dynamical system by discretizing time and iteratively applying
+$\mathbf{x}(t_{s+1})=\mathbf{x}(t_{s})+\Delta t\cdot h(\mathbf{x}(t_{s})),$
+where $h$ denotes the function governing the dynamics of the system.
+In our setting, this transition is modeled as a data-driven one-step mapping $x(t_{s})\to x(t_{s+1})$, learned using GP emulators.
+Importantly, unlike previously established findings (Girard et al.,, 2002; Mohammadi et al.,, 2019) that rely on MC approximation, our method analytically derives the predictive mean and variance using linked GPs, enabling exact and sampling-free uncertainty propagation at each time step.
+This allows the GP-based flow map to serve as a probabilistic surrogate for the unknown vector field $h$, yielding a data-driven and uncertainty-aware generalization of the Euler scheme.
+Building on this foundation, we integrate the concept of linked GP emulators to propagate predictions with uncertain inputs over time.
+Specifically, our method recursively links GP outputs to emulate each state variable, preserving accuracy while maintaining computational efficiency.
+The feed-forward architecture of the proposed framework, illustrated in Figure 1, forms a sequentially layered structure where each predicted state is linked to the previous one.
+Interestingly, this structure can be interpreted as a special case within the emerging field of deep GP (DGP) models (Damianou and Lawrence,, 2013), which have gained significant attention in the computer experiments literature for their hierarchical representation and enhanced flexibility (Ko and Kim,, 2022; Sauer et al.,, 2023; Ming et al.,, 2023).
+However, unlike canonical DGPs that involve latent intermediate layers and require approximate inference, our approach replaces those latent layers with a series of GP emulators, linked recursively to propagate states in a feed-forward manner reminiscent of deep learning models.
 […]
 Figure 1: An illustration of the self-linked GP emulator for a dynamic system governed by $d$ ODEs.
-Starting with the initial state $\mathbf{x}(t_{0})$, each emulator $\hat{f}_{m}$, $m=1,\ldots,d$ predicts the $m$-th component of the state $\hat{x}_{m}(t_{s})$ for $s=1,\ldots,T$.
-The predicted state $\hat{\mathbf{x}}(t_{s})$ is then fed forward as the input for the next time step $s+1$.
-3.1 Algorithm without forcing inputs
-For a nonlinear dynamic simulator $f$ governed by $d$ ODEs, consider GP emulators $\hat{f}_{m}$, $m=1,\ldots,d$ constructed by using an initial design $\mathbf{X}=(\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{n}(t_{0}))$, and corresponding outputs $\mathbf{y}_{m}=(x_{m}^{1}(t_{1}),\ldots,x_{m}^{n}(t_{1}))^{T}$ following the one-step-ahead approach.
-Instead of relying on MC approximations to compute (3) and (2.3), we derive analytical expressions for $\mathbb{E}\left[\hat{x}_{m}(t_{s})\right]$ and $\mathbb{V}\left[\hat{x}_{m}(t_{s})\right]$ under the squared exponential kernel given $\hat{\mathbf{x}}(t_{s-1})$ are as follows:
-$\mu_{m}(\hat{\mathbf{x}}(t_{s}))$ $=\mathbb{E}\left[\hat{x}_{m}(t_{s})\right]=\mathbb{E}\left[\hat{f}_{m}(\hat{\mathbf{x}}(t_{s-1}))\right]$
-$=\begin{pmatrix}1&\hat{\mathbf{x}}(t_{s-1})\end{pmatrix}\beta_{m}+\sum^{n}_{i=1}r_{i}\prod_{l=1}^{d}\sqrt{\frac{\theta_{ml}}{\theta_{ml}+2\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}}\exp{\left(-\frac{\left(x^{i}_{l}(t_{0})-\mu_{l}(\hat{\mathbf{x}}(t_{s-1}))\right)^{2}}{\theta_{ml}+2\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}\right)},$ (5)
+Starting with the initial state $\mathbf{x}(t_{0})$, each emulator $\hat{f}_{m}$, $m=1,\ldots,d$ predicts the $m$-th component of the state $x_{m}(t_{s})$ for $s=1,\ldots,T$.
+The predicted state $\mathbf{x}(t_{s})$ is then fed forward as the input for the next time step $s+1$.
+We now detail the proposed exact and sampling-free, recursive emulation procedure for the dynamic simulator $f$ introduced in Section 2.3.
+As specified in equation (2), the simulator evolves via the one-step transition $\mathbf{x}(t_{s+1})=(\hat{f}_{1}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s})),\ldots,\hat{f}_{d}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s})))^{\top}$, where the forcing input $\mathbf{w}(t_{s})$ may be omitted in autonomous systems.
+To emulate this mapping, we construct $d$ independent GP models $\hat{f}_{m}$, one for each state variable $m=1,\ldots,d$, using the one-step-ahead responses $\mathbf{y}_{m}=(x_{m}^{1}(t_{1}),\ldots,x_{m}^{n}(t_{1}))$.
+Depending on the presence of forcing inputs, the emulator takes the form either $x_{m}(t_{s+1})=\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))$ or $x_{m}(t_{s+1})=\hat{f}_{m}(\mathbf{x}(t_{s}))$.
+Predictions are obtained recursively using the following closed-form expressions for the posterior predictive mean in (3) and variance in (2.3) under a squared exponential kernel, originally derived by Kyzyurova et al., (2018) and adapted here for the one-step-ahead emulation setting:
+$\mathbb{E}(x_{m}(t_{s+1}))=\begin{pmatrix}1&\mathbf{x}(t_{s})\end{pmatrix}\beta_{m}$
+$\quad+\sum^{n}_{i=1}r_{i}\prod_{j=1}^{p}\exp\left(-\frac{\left(w^{i}_{j}(t_{0})-w_{j}(t_{{s+1}})\right)^{2}}{\theta_{wj}}\right)\prod_{l=1}^{d}\sqrt{\frac{\theta_{ml}}{\theta_{ml}+2\mathbb{V}(x_{l}(t_{s}))}}\exp{\left(-\frac{\left(x^{i}_{l}(t_{0})-\mathbb{E}(x_{l}(t_{s}))\right)^{2}}{\theta_{ml}+2\mathbb{V}(x_{l}(t_{s}))}\right)},$ (5)
 and
-$\sigma^{2}_{m}(\hat{\mathbf{x}}(t_{s}))$ $=\mathbb{V}\left[\hat{x}_{m}(t_{s})\right]=\mathbb{V}\left[\hat{f}_{m}(\hat{\mathbf{x}}(t_{s-1}))\right]$
-$=\tau^{2}_{m}-\left(\mu_{m}(\hat{\mathbf{x}}(t_{s}))-\begin{pmatrix}1&\hat{\mathbf{x}}(t_{s-1})\end{pmatrix}\beta_{m}\right)^{2}+\left(\sum_{i,k=1}^{n}\left(r_{i}r_{k}-\tau^{2}_{m}(\mathbf{K}^{-1}_{m})_{ik}\right)\prod_{l=1}^{d}\zeta_{lik}\right),$ (6)
-where $r_{i}=(\mathbf{K}^{-1}_{m}(\mathbf{y}_{m}-\mathbf{X}\beta_{m}))_{i}$, and
-$\zeta_{lik}=\sqrt{\frac{\theta_{ml}}{\theta_{ml}+4\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}}\exp{\left(-\frac{\left(\frac{x^{i}_{l}(t_{0})+x^{k}_{l}(t_{0})}{2}-\mu_{l}(\hat{\mathbf{x}}(t_{s-1}))\right)^{2}}{\frac{\theta_{ml}}{2}+2\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}-\frac{(x^{i}_{l}(t_{0})-x^{k}_{l}(t_{0}))^{2}}{2\theta_{ml}}\right)}.$
-The closed-form expressions for the posterior predictive mean (3) and variance (2.3) under a Matérn kernel with smoothness parameters $\nu=1.5$ and $\nu=2.5$ can be straightforwardly derived following the proof in Ming and Guillas, (2021).
-Consistent with prior studies (Girard et al.,, 2002; Kyzyurova et al.,, 2018; Mohammadi et al.,, 2019; Ming and Guillas,, 2021), we adopt the moment matching method to approximate the posterior distribution as Gaussian, leveraging the mean and variance from (5) and (6).
-The moment matching approximation suits the structure of self-linked GP emulator, where outputs recursively serve as inputs, since the method aligns with the inherent Gaussian nature of GPs and ensures analytical tractability and recursive consistency.
-Algorithm 1 Exact emulation of dynamic simulators without forcing inputs
+$\mathbb{V}(x_{m}(t_{s+1}))=\tau^{2}_{m}-\left(\mathbb{E}(x_{m}(t_{s+1}))-\begin{pmatrix}1&\mathbf{x}(t_{s})\end{pmatrix}\beta_{m}\right)^{2}$
+$\quad+\sum_{i,k=1}^{n}\left(r_{i}r_{k}-\tau^{2}_{m}(\mathbf{K}^{-1}_{m})_{ik}\right)\prod_{j=1}^{p}\exp{\left(-\frac{\left(w^{i}_{j}(t_{0})-w_{j}(t_{s+1})\right)^{2}+\left(w^{k}_{j}(t_{0})-w_{j}(t_{s+1})\right)^{2}}{\theta_{wj}}\right)}\prod_{l=1}^{d}\zeta_{lik},$ (6)
+where $r_{i}=(\mathbf{K}^{-1}_{m}(\mathbf{y}_{m}-\begin{pmatrix}1&\mathbf{X}\end{pmatrix}\beta_{m}))_{i}$ and
+$\zeta_{lik}=\sqrt{\frac{\theta_{ml}}{\theta_{ml}+4\mathbb{V}(x_{l}(t_{s}))}}\exp{\left(-\frac{\left(\frac{x^{i}_{l}(t_{0})+x^{k}_{l}(t_{0})}{2}-\mathbb{E}(x_{l}(t_{s}))\right)^{2}}{\frac{\theta_{ml}}{2}+2\mathbb{V}(x_{l}(t_{s}))}-\frac{(x^{i}_{l}(t_{0})-x^{k}_{l}(t_{0}))^{2}}{2\theta_{ml}}\right)}.$
+Here, $\theta_{ml}$ and $\theta_{wj}$ are the lengthscale hyperparameters associated with the state and forcing input dimensions, respectively.
+Algorithm 1 Exact emulation of dynamic simulators
 1:for $m=1$ to $d$ do
-2:     Fit $\hat{f}_{m}$ using $\mathbf{X}=(\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{n}(t_{0}))$, and $\mathbf{y}_{m}=(x^{1}_{m}(t_{1}),\ldots,x^{n}_{m}(t_{1}))^{T}$
+2:  Fit $\hat{f}_{m}$ on training data $(\mathbf{X},\mathbf{W},\mathbf{y}_{m})$
 3:end for
-4:Set $\hat{\mathbf{x}}(t_{0})\leftarrow\mathbf{x}(t_{0})$
-5:for $s=0$ to $T-1$ do
+4:Set $\mathbf{x}(t_{0})\leftarrow\mathbf{x}(t_{0})$
+5:for $s=0$ to T1T-1 do
 6:  for $m=1$ to $d$ do
-7:         Compute $\mu_{m}(\hat{\mathbf{x}}(t_{s+1}))$ using (5)
-8:         Compute $\sigma^{2}_{m}(\hat{\mathbf{x}}(t_{s+1}))$ using (6)
+7:   Compute $\mathbb{E}(x_{m}(t_{s+1}))$ using (5)
+8:   Compute $\mathbb{V}(x_{m}(t_{s+1}))$ using (6)
 9:  end for
@@
 11:end for
-12:Return: $\hat{\mathbf{x}}(t_{s})$ for $s=1,\ldots,T$
-By leveraging these closed-form expressions, computationally expensive MC approximations can be avoided, allowing for exact computations.
-The elimination of MC sampling significantly reduces computational costs, making the proposed approach particularly suitable for dynamical systems with many state variables that require extensive simulations.
-Moreover, the proposed framework provides exact predictive moments, ensuring both efficiency and accuracy.
-This is particularly advantageous for long-term iterative predictions, where even small numerical errors in early steps can accumulate over time, leading to significant deviations in later steps—a phenomenon often referred to as the butterfly effect.
-By directly computing the mean and variance, our approach mitigates this issue, enhancing robustness in sequential simulations.
-Furthermore, multi-step-ahead predictions can be efficiently obtained by recursively applying (5) and (6), reducing the need for costly recomputation at each iteration.
-3.2 Algorithm with forcing input
-Assume we have a known, $d_{w}$-dimensional forcing input $\mathbf{w}(t_{s})$ at time $t_{s}$ for $s=1,\ldots,T$. The dynamic simulator $f$ can be expressed as:
-$x_{m}(t_{s+1})=f_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s})),\quad\text{for}\quad m=1,\ldots,d.$ (7)
-Accordingly, the Markov property must be adjusted to account for forcing inputs. Instead of (2), the emulator follows:
-$\hat{x}_{m}(t_{s+1})=\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s})),$ (8)
-where $m=1,\ldots,d$.
-The algorithm proposed in Section 3.1 can be naturally extended to incorporate forcing inputs by treating them as additional variables in the GP framework.
-This extension preserves the computational efficiency of the method while enabling accurate predictions for systems influenced by external inputs.
-To account for forcing inputs, we modify the analytical expressions for $\mathbb{E}\left[\hat{x}_{m}(t_{s})\right]$ and $\mathbb{V}\left[\hat{x}_{m}(t_{s})\right]$ under the squared exponential kernel given $\hat{\mathbf{x}}(t_{s-1})$ are as follows:
-$\mu_{m}(\hat{\mathbf{x}}(t_{s}))=\begin{pmatrix}1&\hat{\mathbf{x}}(t_{s-1})\end{pmatrix}\beta_{m}$
-$+\sum^{n}_{i=1}r_{i}\prod_{j=1}^{d_{w}}\exp\left(-\frac{\left(w^{i}_{j}(t_{0})-w_{j}(t_{s})\right)^{2}}{\theta_{wj}}\right)\prod_{l=1}^{d}\frac{1}{\sqrt{1+2\frac{\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}{\theta_{ml}}}}\exp{\left(-\frac{\left(x^{i}_{l}(t_{0})-\mu_{l}(\hat{\mathbf{x}}(t_{s-1}))\right)^{2}}{\theta_{ml}+2\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}\right)},$ (9)
-and
-$\sigma^{2}_{m}(\hat{\mathbf{x}}(t_{s}))=\tau^{2}_{m}-\left(\mu_{m}(\hat{\mathbf{x}}(t_{s}))-\begin{pmatrix}1&\hat{\mathbf{x}}(t_{s-1})\end{pmatrix}\beta_{m}\right)^{2}$
-$+\left(\sum_{i,k=1}^{n}\left(r_{i}r_{k}-\tau^{2}_{m}(\mathbf{K}^{-1}_{m})_{ik}\right)\prod_{j=1}^{d_{w}}\exp{\left(-\frac{\left(w^{i}_{j}(t_{0})-(w_{j}(t_{s})\right)^{2}+\left(w^{k}_{j}(t_{0})-w_{j}(t_{s})\right)^{2}}{\theta_{wj}}\right)}\prod_{l=1}^{d}\zeta_{lik}\right),$ (10)
-where $\theta_{wj}$ for $j=1,\ldots,d_{w}$ are the lengthscale hyperparameters associated with the forcing inputs.
-Algorithm 2 Exact emulation of dynamic simulators with forcing inputs
-1:for $m=1$ to $d$ do
-2:     Fit $\hat{f}_{m}$ using $\mathbf{X}=(\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{n}(t_{0}))$, $\mathbf{W}=(\mathbf{w}^{1}(t_{0}),\ldots,\mathbf{w}^{n}(t_{0}))$, and $\mathbf{y}_{m}=(x^{1}_{m}(t_{1}),\ldots,x^{n}_{m}(t_{1}))^{T}$
-3:end for
-4:Set $\hat{\mathbf{x}}(t_{0})\leftarrow\mathbf{x}(t_{0})$
-5:for $s=0$ to $T-1$ do
-6:     for $m=1$ to $d$ do
-7:         Compute $\mu_{m}(\hat{\mathbf{x}}(t_{s+1}))$ using (9) with $\mathbf{w}(t_{s})$
-8:         Compute $\sigma^{2}_{m}(\hat{\mathbf{x}}(t_{s+1}))$ using (10) with $\mathbf{w}(t_{s})$
-9:     end for
-10:     Update $s\leftarrow s+1$
-11:end for
-12:Return: $\hat{\mathbf{x}}(t_{s})$ for $s=1,\ldots,T$
-The additional term in these expressions corresponds to the kernel function $K(\mathbf{w}(t_{0}),\mathbf{w}(t_{s}))$, which arises because the forcing inputs $\mathbf{w}(t_{s})$ are treated as known rather than random.
-This assumption simplifies the model structure, as the forcing inputs remain unlinked variables in the emulator.
+12:Return: $\mathbf{x}(t_{s})$ for $s=1,\ldots,T$
+The forcing inputs are treated as known and contribute deterministically through their kernel terms, without introducing additional uncertainty.
+When forcing inputs are absent, the expressions simplify by setting $p=0$.
+The closed-form expressions for the posterior predictive mean and variance under a Matérn kernel can be straightforwardly derived following the development in Ming and Guillas, (2021).
+Consistent with prior studies (Girard et al.,, 2002; Kyzyurova et al.,, 2018; Mohammadi et al.,, 2019; Ming and Guillas,, 2021), we adopt the moment matching method to approximate the posterior distribution as Gaussian, leveraging the analytic mean and variance in (5) and (6).
+This approach is particularly well-suited to the self-coupled structure of the proposed framework, where the prediction at each step serve as input to the next time step.
+By computing the mean and variance analytically, the proposed sampling-free method achieves exact moment propagation, avoiding the accumulation of sampling error.
+This is especially beneficial for predictions over long time horizon, where even small numerical inaccuracies from early steps can compound over time, leading to significant degradations in later stages—a phenomenon commonly referred to as the butterfly effect.
+Moreover, the feed-forward structure supports efficient multi-step-ahead predictions without repeated recomputation of previous states, providing further computational gains.
 4 Numerical Studies
 In this section, we conduct a series of numerical experiments to evaluate the performance of the proposed approach.
-Specifically, we compare the predictive performance of the proposed exact algorithm with an emulation method based on MC approximation, using $n_{MC}=1000$ samples.
-Both approaches employ the same GP emulator, using an anisotropic squared exponential kernel and a linear mean function $\alpha(\mathbf{x})=(1,\mathbf{x}^{T})\beta$.
-All numerical studies are conducted using a maximin Latin hypercube design (Johnson et al.,, 1990; Morris and Mitchell,, 1995), with the number of initial samples set to $n=12d$, following the guidelines of Mohammadi et al., (2019), which builds on the design principles established in Jones et al., (1998) and Loeppky et al., (2009).
+Specifically, we compare the predictive performance of the proposed exact algorithm with an emulation method based on MC approximation, considering three sample sizes $(n_{MC}=10,100,1000)$.
+Importantly, both approaches adopt the moment-matching method, and employ the same underlying GP emulator, constructed using an anisotropic squared exponential kernel and a linear mean function $\alpha(\mathbf{x})=(1,\mathbf{x})\beta$.
+The difference between the two methods lies solely in how predictive moments are computed as described in Section 2: the MC-based approach relies on sampling, whereas the proposed exact method computes them analytically.
+All numerical experiments are conducted using a maximin Latin hypercube design (Johnson et al.,, 1990; Morris and Mitchell,, 1995), with the number of initial samples set to $n=10\times d$, following the guidelines of Jones et al., (1998) and Loeppky et al., (2009).
 The time step difference is set to $\Delta t=t_{s}-t_{s-1}=0.01$.
@@
 This solver, known as the Livermore solver for stiff and non-stiff ODE systems, automatically switches between stiff and non-stiff methods for efficiency (Petzold,, 1983).
-To evaluate predictive performance, we compute predictions at all time points $t_{s}$, $s=1,\ldots,T$, and assess model accuracy using the root-mean-square error (RMSE) and continuous rank probability score (CRPS) (Gneiting and Raftery,, 2007) over 100 repetitions.
+Predictive performance is evaluated across all time points $t_{s}$, $s=1,\ldots,T$, using two standard metrics: the root-mean-square error (RMSE) and the mean absolute error (MAE), averaged over 100 repetitions.
 The evaluation metrics are defined as follows:
-RMSE $=\sqrt{\sum_{i=1}^{n}\frac{\left(y_{i}-\hat{y}_{i}\right)^{2}}{n}},$
-CRPS $=\sigma\left[\frac{1}{\sqrt{\pi}}-2\phi\left(\frac{y-\mu}{\sigma}\right)-\frac{y-\mu}{\sigma}\left(2\Phi\left(\frac{y-\mu}{\sigma}\right)-1\right)\right],$
-where $y$ represents the true function values, $\mu$ and $\sigma^{2}$ denote the posterior predictive mean and variance, respectively.
-$\phi$ and $\Phi$ are the pdf and the cdf of a standard Gaussian distribution.
-These metrics quantify the degree of discrepancy between the true function and the predictive model, with lower RMSE and CRPS values indicating better model accuracy.
-Furthermore, we evaluate computational efficiency by comparing computation times across different emulation approaches.
+RMSE $=\sqrt{\frac{1}{T}\sum_{s=1}^{T}\left(x_{m}(t_{s})-\mathbb{E}(x_{m}(t_{s}))\right)^{2}},$
+MAE $=\frac{1}{T}\sum_{s=1}^{T}|x_{m}(t_{s})-\mathbb{E}(x_{m}(t_{s}))|.$
+These metrics measure the discrepancy between the true function and the predictive model, with lower RMSE and MAE values indicating higher accuracy.
+Furthermore, computational efficiency is assessed by comparing computation times across different emulation approaches.
+In addition to prediction accuracy, we assess how predictive uncertainty accumulates over time.
+Specifically, we determine the number of steps for which predictions remain reliable and the point at which uncertainty increases substantially.
+For this purpose, we apply change point detection (Killick et al.,, 2012) to the predictive standard deviation using the cpt.mean function in the changepoint package (Killick and Eckley,, 2014).
+The comparison is performed only between the exact method and the MC approach with 1000 samples, as the latter provides the most accurate MC-based predictions and yields more stable results by reducing variability from sampling.
+This analysis addresses the practical question of how far ahead the emulator can provide useful predictions before the uncertainty becomes large enough to degrade performance.
 4.1 Lotka-Volterra model
@@
 For this study, we consider a simulation over $t\in[0,30]$, with prespecified parameter values $r_{G}=1.5$, $K=10$, $r_{I}=1.5$, $k_{AE}=1$, and $r_{M}=2$.
-Initial samples are drawn from $[0,5]^{2}$.
+Initial samples are drawn from $0,5]^{2}$.
 The objective is to predict the system’s dynamics given the initial state $\mathbf{x}(t_{0})=(P(t_{0})=1,C(t_{0})=2)$.
-Figure 2: RMSE (first and second plots), CRPS (third and fourth plots) of emulation algorithms to $P$ and $C$ and computation time (fifth plots) across 100 repetitions for Lotka-Volterra model.
-Figure 3: Lotka-Volterra model (black solid line) and emulation results of prediction (green dashed line) via each algorithm.
-Figure 2 presents the emulation results over 100 repetitions for RMSE and CRPS, while Figure 3 illustrates the predicted trajectories for state variables, $P$ and $C$, respectively.
-As expected, both algorithms produce nearly identical predictions, with a maximum difference of only 0.000024.
-While both methods achieve high predictive accuracy, their computational costs differ considerably.
-On average, the exact algorithm requires only 0.40603 seconds per repetition, whereas the MC approach takes 90.13912 seconds.
+Figure 2: RMSE (first and second panels), MAE (third and fourth panels) of emulation algorithms to $P$ and $C$ and computation time (fifth panel) across 100 repetitions for the Lotka-Volterra model.
+Figure 2 presents the emulation results over 100 repetitions for RMSE and MAE.
+As expected, the MC approach produces suboptimal performance, although larger sample sizes yield results closer to those of the exact method.
+Compared with the exact method, the MC approach with 10 samples yields RMSE and MAE values that are 6–8% larger, with 100 samples yields values that are 4–6% larger, and with 1000 samples yields values that are less than 1% larger.
+Moreover, their computational costs differ considerably.
+On average, the exact algorithm requires only 0.45031 seconds per repetition, whereas the MC approach takes 0.48785 seconds with 10 samples, 1.70622 seconds with 100 samples, and 101.37841 seconds with 1000 samples.
 This comparison highlights the advantage of the proposed framework in significantly reducing computational cost while maintaining accuracy in dynamic system emulation.
-Moreover, the computational gap would increase further if the MC approach employed a larger number of samples, or if the simulation involved many time steps or a longer time horizon.
-Figure 4: Boxplots of computation time for the Exact and Monte Carlo methods as the number of design points varies from 5 to 50 in increments of 5 across 100 repetitions for Lotka-Volterra model.
-Figure 4 illustrates the computation time for emulation as the number of samples varies from 5 to 50 per repetition.
-With a precomputed $\mathbf{K}^{-1}$, the computational complexity of computing the predictive mean and variance at a single input location using MC-based emulation is $\mathcal{O}(n^{2}n_{MC})$.
-On the other hand, under the same conditions, the exact algorithm requires only $\mathcal{O}(n^{2}d)$ complexity.
-Since $n_{MC}$ is typically in the range of several thousand or more, the proposed algorithm offers substantial computational savings.
-Figure 5: RMSE (first plot), CRPS (second plot) of emulation algorithms to $P$ with forcing input $C$ and computation time (third plot) across 100 repetitions for Lotka-Volterra model.
-Figure 5 demonstrates the results for the case in which $C(t)$ is treated as a forcing input.
-As anticipated, the emulation performance improves due to the availability of a known input, and the computation time is halved with 0.20403 seconds per repetition for the exact algorithm and 48.03827 seconds for the MC-based algorithm.
-These results confirm that the exact method remains computationally efficient in the presence of a forcing input.
+As a result, the computational gap would increase further if the MC approach employed a larger number of samples, or if the simulation involved many time steps or a longer time horizon.
+Figure 3: Lotka-Volterra model (black solid line) and emulation results of prediction (green dashed line) via each algorithm with detected change points (red dotted vertical line)
+Figure 3 illustrates the predicted trajectories for state variables, $P$ and $C$, along with the change point of predictive uncertainty.
+Because the one-step-ahead approach is effective enough for emulating the Lotka-Volterra model, both methods accurately predict the simulator outputs with very small uncertainties of approximately $10^{-4}$ for the exact method and $10^{-2}$ for MC method.
+The Exact method does not detect any change points within the given time steps, whereas the MC method detects changes in uncertainties at $t=5.53$ for $P$ and $t=3.33$ for $C$.
+4.2 Duffing oscillator
+The Duffing oscillator (Duffing,, 1918; Ueda,, 1991) is a canonical model in nonlinear dynamics that describes the behavior of a damped oscillator subject to a nonlinear restoring force.
+It captures the dynamics of a second-order system defined in terms of the displacement variable $x(t)$ and its velocity $v(t)$, governed by the following system of differential equations:
+$\begin{cases}\frac{dx}{dt}&=v,\\ \frac{dv}{dt}&=-\delta v-\alpha x-\beta x^{3},\end{cases}$
+where $\delta>0$ is the damping coefficient, $\alpha$ controls the linear stiffness, and $\beta$ governs the strength of the nonlinear restoring force.
+In this study, we simulate the system over $t\in[0,30]$ with parameter values $\delta=0.1$, $\alpha=1$, and $\beta=1$.
+The initial state is set as $\mathbf{x}(t_{0})=(x(t_{0})=1,v(t_{0})=0)$, and training inputs are sampled from the domain $1,1]^{2}$.
+Figure 4: RMSE (first and second panels), MAE (third and fourth panels) of emulation algorithms to $x$ and $v$ and computation time (fifth panel) across 100 repetitions for the Duffing oscillator.
+Figure 4 illustrates the Duffing oscillator results, which exhibit patterns similar to those observed for the Lotka–Volterra model.
+The MC approach again underperforms relative to the exact method, although the performance gap narrows as the number of samples increases.
+With 10 MC samples, RMSE and MAE are approximately 24–25% larger than those of the exact method; with 100 samples, the difference drops to 5–6%, and with 1000 samples, to about 2.5%.
+Computational times show the same trend as before, with the exact algorithm averaging 0.52205 seconds per repetition compared to 0.47833, 1.69159, and 107.25105 seconds for the MC approach with 10, 100, and 1000 samples, respectively.
+Figure 5: Duffing oscillator (black solid line) and emulation results of prediction (green dashed line) via each algorithm with detected change points (red dotted vertical line)
+Again, the change point detection results for the Duffing oscillator in Figure 5 follow the same pattern observed for the Lotka–Volterra model.
+Both methods yield accurate predictions, with predictive uncertainties around $10^{-6}$ for the exact method and $10^{-4}$ for the MC method.
+As in the previous case, the exact method does not detect any change points, whereas the MC method identifies changes at $t=12.35$ for $x$ and $t=5.51$ for $v$.
 4.3 Lorenz system
@@
 where $X$, $Y$, and $Z$ represent the system’s state variables, and $a$, $b$, and $c$ are system parameters that govern the system’s dynamics.
-Using the classic values $a=-8/3$, $b=-10$, and $c=28$, originally employed by Lorenz to study convection rolls, the system exhibits chaotic behavior.
+Using the classic values a=8/3a=-8/3, b=10b=-10, and $c=28$, originally employed by Lorenz to study convection rolls, the system exhibits chaotic behavior.
 We focus on a simulation over $t\in[0,25]$, and the initial state $\mathbf{x}(t_{0})=(X(t_{0})=-1,Y(t_{0})=-1,Z(t_{0})=-1)$.
-Initial designs are sampled from $[-10,10]^{3}$.
-Figure 6: RMSE (first to third plots), CRPS (fourth to sixth plots) of emulation algorithms to $P$ and $C$ and computation time (seventh plots) across 100 repetitions for Lorenz system.
-Figure 7: Lorenz system (black solid line) and emulation results of prediction (green dashed line) via each algorithm.
-Figure 8: 3-d illustrations of Lorenz system (black solid line) and emulation results of prediction (green dashed line) via each algorithm.
-From Figure 6, the emulation results show slight differences between the two algorithms.
-This discrepancy arises because the Lorenz system is highly sensitive to small changes, which can lead to significant differences, a phenomenon known as the “butterfly effect”.
-As shown, the exact algorithm can emulate the system efficiently, requiring only 0.95777 seconds per repetition, whereas the MC approach takes significantly longer at 143.22403 seconds.
-Figures 7 and 8 illustrate the predicted trajectories for each state variable, $X$, $Y$, and $Z$.
-Both algorithms provide accurate predictions at early time steps up to $t=20$, but their performance deteriorates beyond this point.
+Initial designs are sampled from $10,10]^{3}$.
+Figure 6: RMSE (first to third panels), MAE (fourth to sixth panels) of emulation algorithms to $X$, $Y$ and $Z$ and computation time (seventh panel) across 100 repetitions for Lorenz system.
+Figure 6 shows that the Lorenz system results differ notably from those of the previous two simulators.
+The Lorenz system is a more sensitive and complex dynamical system, where small changes can lead to large deviations, a phenomenon known as the “butterfly effect”.
+This greater sensitivity makes the relative differences between the exact method and MC approaches more pronounced at small sample sizes and more inconsistent across state variables.
+With 10 MC samples, RMSE and MAE are 6–17% larger than those of the exact method, with 100 samples, the difference decreases to 1.5–4%, and with 1000 samples, it ranges from 0% to 3.5%.
+Computational costs remain substantially different, with the exact algorithm requiring 0.99940 seconds per repetition compared to 0.65037, 2.69461, and 160.60290 seconds for the MC approach with 10, 100, and 1000 samples, respectively.
+These results highlight that the MC approach requires a sufficiently large number of samples to achieve performance comparable to the exact method, but this comes at a substantial increase in computational time.
+Figure 7: Lorenz (black solid line) and emulation results of prediction (green dashed line) via each algorithm with detected change points (red dotted vertical line) and corresponding 3-D trajectory illustrations.
+Figure 7 illustrates the predicted trajectories with change point detection for each state variable, $X$, $Y$, and $Z$.
+Both algorithms produce accurate predictions at early time steps up to $t=20$, but performance deteriorates thereafter.
+For the exact method, the first time step at which the absolute difference between the prediction and the true output exceeds 10 is $t=19.88$ for $X$, $t=20.69$ for $Y$, and $t=19.99$ for $Z$, which is later than the corresponding times for the MC method, which are $t=19.77$ for $X$, $t=20.47$ for $Y$, and $t=19.90$ for $Z$.
 For highly sensitive dynamic systems such as the Lorenz system, the exact algorithm is particularly advantageous, as MC methods introduce variations that exacerbate prediction errors over time.
-This underscores the importance of exact emulation for improving accuracy and computational efficiency in chaotic systems.
+This is consistent with the change point detection results, where the MC approach identifies a change point at $t=0.12$ for $X$ which is near the very beginning of the time steps, and at $t=19.29$ for $Y$ and $t=19.27$ for $Z$.
+In contrast, exact method maintains smaller uncertainties and detects only a single change point at $t=0.3$ for $X$ caused by increased uncertainty at the peak behavior, while no change points are detected for $Y$ and $Z$ within the given time steps under the same penalty value in the change point detection analysis.
 5 Discussion
 In this paper, we propose fast and exact algorithms for emulating complex dynamic simulators.
-These algorithms enable efficient emulation compared to traditional MC-based approaches, while retaining their advantages.
-Numerical studies demonstrate that our method achieves equivalent results with significantly reduced computation time, over 100 times faster.
-Additionally, our approach accurately computes the predictive posterior mean and variance, preventing the accumulation of small differences from sampling errors.
-From a computational perspective, the addition of jitter for numerical stability–commonly used to avoid ill-conditioned covariance matrices–may introduce minor discrepancies that could accumulate over time.
+These algorithms offer substantial improvements over traditional MC-based approaches by retaining their key advantages while significantly reducing computational overhead.
+Specifically, the proposed method bypasses MC approximations by deriving closed-form expressions for the posterior predictive mean and variance.
+This not only yields more than 150-fold improvements in computational efficiency without compromising accuracy, but also prevents the accumulation of small differences from sampling errors, thereby ensuring stable propagation over time.
+The framework also naturally accommodates forcing inputs by treating them as additional input variables within the GP formulation, enabling emulation of a broader class of dynamic systems.
+While uncertainty quantification is also supported in MC-based methods, our method preserves this capability and provides predictive variance analytically at each time step, supporting reliable uncertainty-aware emulation.
+To facilitate adoption and reproducibility, we provide an R package, dynemu, available on CRAN, which implements the proposed methodology in a user-friendly and computationally efficient framework.
+From a computational perspective, the addition of jitter for numerical stability (Ranjan et al.,, 2011)–commonly used to avoid ill-conditioned covariance matrices–may introduce minor discrepancies that could accumulate over time.
 In dynamic systems that are highly sensitive to small changes, such as those exhibiting the butterfly effect, these numerical artifacts can significantly impact the emulation outcomes.